# Limits of Probability Measures with General Coefficients

### v2

@@ 6.2. Application to the $\beta$-Hermite ensemble
We discuss an example of computing the coefficient of a term, where the coefficient satisfies the conditions of Lemma 6.4.
@@
Suppose $\mathfrak{r}$ consists of $2k$ equal positive integers and $\theta=1$. For $N\geq\max(\mathfrak{r})$, the coefficient of $(c_{F}^{(2)})^{k}$ in $[1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))$ is
@@
see Proposition 4.4. Then, using the calculation of $\mathbb{E}_{\mu_{N}}\left[\sum_{j=1}^{N}a_{j}^{2k}\right]$ in [eulerchar] finishes the proof. ∎
@@
Similarly, using [random_matrix]*(6.5.30) gives an expression for the coefficient of $(c_{F}^{(2)})^{k}$ in $[1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))$ for general $\mathfrak{r}$ when $\theta=1$.
@@
Suppose $\mathfrak{r}$ consists of $2k$ equal positive integers. For $N\geq\text{max}(\mathfrak{r})$, the coefficient of $(c_{F}^{(2)})^{k}$ in $[1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))$ is
@@
First, observe that the coefficient of $N^{k}$ is a multiple of $\theta^{k}$ because the only method to contribute a factor of $N^{k}$ is to have $k$ switches and $k$ term multiplications, see Lemma 6.4. By setting $\theta=1$ and using Lemma 6.7, we obtain the coefficient of $N^{k}$.
@@
Without loss of generality, assume that $\mathfrak{r}$ consists of $2k$ $1$s. The coefficient of $\theta^{k-1}N^{k-1}$ is contributed to by sequences of $k$ term multiplications, $k-1$ switches, and $1$ derivative.
@@
Furthermore, in order to achieve the factor $N^{k-1}$, each of the $k-1$ switches are a switch from $1$ to a distinct index $i\in[N]\backslash\{1\}$. If some of the indices are the same, then we would contribute a factor of at most $N^{k-2}$. The number of ways to choose the indices of the switches in this case is $(k-1)!\binom{N-1}{k-1}$; since we are only considering the coefficient of $\theta^{k-1}N^{k-1}$, the number of ways to choose the indices is equivalent to $N^{k-1}$. By symmetry, we can assume that the indices from the first switch to the last are $2,3,\ldots,k$.
@@
Note that this argument is from [matrix], which uses it to justify replacing the switch from $i_{j}$ to $i$ with $d_{i_{j}}$ in step 3 of the proof of Proposition 5.5. We similarly replace the switch from $1$ to $i$ with $d_{1}$, which contrasts with the proof of 3.17.2 where we would replace the switch with the change from $1$ to $i$, see Remark 3.18.
Let $\zeta$ by some ordering of the $k$ term multiplications by $2x_{1}$, $k-1$ switches, and $1$ derivative $\partial_{1}$. The value of $r(s)_{2k}$ for each of the $N^{k-1}+O(N^{k-2})$ sequences that are equivalent to $\zeta$ is the same; let this common value be $v(\zeta)\theta^{k-1}$, where the factor of $\theta^{k-1}$ is from the $k-1$ switches. Then, the coefficient of $N^{k-1}\theta^{k-1}$ is $\sum_{\zeta}v(\zeta)$.
@@
Note that the coefficient of $N^{k-1}$ of $(c_{F}^{(2)})^{k}$ is 00 when $\theta=1$, see Lemma 6.7. This implies that the coefficient of $N^{k-1}\theta^{k}$ is the coefficient of $N^{k-1}\theta^{k-1}$ times negative one. ∎
As mentioned earlier, the leading order term of the $2k$th moment of the $\beta$-Hermite ensemble $d_{N,\beta}$ is known to be $\frac{N^{k}}{k+1}\binom{2k}{k}$, see Proposition 5.3 or [eigenvalstat]*Theorem 6.2.3. As an example of an application of this paper’s results, we approximate the lower order terms. Particularly, using Lemma 6.9, we can estimate the moments of the $\beta$-Hermite ensemble for all $\beta>0$ with error $O(N^{-2})$, where $\beta=2\theta$. Note that for the cases $\beta\in\{1,2,4\}$, we can arrive at these estimates using recursion, see [eulerchar] for the GUE case and [momentsrecursion] for the GOE and GSE cases. The following result is true for general $\beta$.
@@
From (22), $c_{F}^{(2)}=\frac{1}{2\theta}$. Then, from Proposition 4.4 and Lemma 6.9, the $2k$th moment is

### v3

@@ 6.2. Application to the $\beta$-Hermite ensemble
We discuss an example of computing the coefficient of a term, where the coefficient satisfies the conditions of Lemma˜6.4.
@@
Suppose $\mathfrak{r}$ consists of $2k$ equal positive integers and $\theta=1$. For $N\geq\max(\mathfrak{r})$, the coefficient of $(c_{F}^{(2)})^{k}$ in $1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))$ is
@@
see Proposition˜4.5. Then, using the calculation of $\mathbb{E}_{\mu_{N}}\left[\sum_{j=1}^{N}a_{j}^{2k}\right]$ in [eulerchar] finishes the proof. ∎
@@
Similarly, using [random_matrix]*(6.5.30) gives an expression for the coefficient of $(c_{F}^{(2)})^{k}$ in $1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))$ for general $\mathfrak{r}$ when $\theta=1$.
@@
Suppose $\mathfrak{r}$ consists of $2k$ equal positive integers. For $N\geq\text{max}(\mathfrak{r})$, the coefficient of $(c_{F}^{(2)})^{k}$ in $1]\mathcal{D}_{\mathfrak{r}}^{\theta}(F(x_{1},\ldots,x_{N}))$ is
@@
First, observe that the coefficient of $N^{k}$ is a multiple of $\theta^{k}$ because the only method to contribute a factor of $N^{k}$ is to have $k$ switches and $k$ term multiplications, see Lemma˜6.4. By setting $\theta=1$ and using Lemma˜6.7, we obtain the coefficient of $N^{k}$.
@@
Without loss of generality, assume that $\mathfrak{r}$ consists of $2k$ 11s. The coefficient of $\theta^{k-1}N^{k-1}$ is contributed to by sequences of $k$ term multiplications, k1k-1 switches, and $1$ derivative.
@@
[…]
@@
Note that this argument is from [matrix], which uses it to justify replacing the switch from $i_{j}$ to $i$ with $d_{i_{j}}$ in step 3 of the proof of Proposition 5.5. We similarly replace the switch from $1$ to $i$ with $d_{1}$, which contrasts with the proof of ˜3.17.2 where we would replace the switch with the change from $1$ to $i$, see Remark˜3.18.
Let $\zeta$ by some ordering of the $k$ term multiplications by $2x_{1}$, k1k-1 switches, and $1$ derivative $\partial_{1}$. The value of $r(s)_{2k}$ for each of the $N^{k-1}+O(N^{k-2})$ sequences that are equivalent to $\zeta$ is the same; let this common value be $v(\zeta)\theta^{k-1}$, where the factor of $\theta^{k-1}$ is from the k1k-1 switches. Then, the coefficient of $N^{k-1}\theta^{k-1}$ is $\sum_{\zeta}v(\zeta)$.
@@
Note that the coefficient of $N^{k-1}$ of $(c_{F}^{(2)})^{k}$ is 0 when $\theta=1$, see Lemma˜6.7. This implies that the coefficient of $N^{k-1}\theta^{k}$ is the coefficient of $N^{k-1}\theta^{k-1}$ times negative one. ∎
As mentioned earlier, the leading order term of the $2k$th moment of the $\beta$-Hermite ensemble $d_{N,\beta}$ is known to be $\frac{N^{k}}{k+1}\binom{2k}{k}$, see Proposition˜5.3 or [eigenvalstat]*Theorem 6.2.3. As an example of an application of this paper’s results, we approximate the lower order terms. Particularly, using Lemma˜6.9, we can estimate the moments of the $\beta$-Hermite ensemble for all $\beta>0$ with error $O(N^{-2})$, where $\beta=2\theta$. Note that for the cases $\beta\in\{1,2,4\}$, we can arrive at these estimates using recursion, see [eulerchar] for the GUE case and [momentsrecursion] for the GOE and GSE cases. The following result is true for general $\beta$.
@@
From (22), $c_{F}^{(2)}=\frac{1}{2\theta}$. Then, from Proposition˜4.5 and Lemma˜6.9, the $2k$th moment is
@@
6.3. Generalization to the $|\theta_{N}N|\rightarrow\infty$ regime
We note that if $\lim_{N\rightarrow\infty}\theta_{N}N^{c}=\theta$ for $\theta\in\mathbb{C}$ nonzero and $c<1$, then $|\theta_{N}N|\rightarrow\infty$. Based on this observation, we consider the regime $|\theta_{N}N|\rightarrow\infty$. This regime is considered for the $\beta$-Hermite ensemble in [betagaussian] and for the $\beta$-Laguerre ensemble in [betalaguerre]. Since the Bessel generating function for the $\beta$-Hermite ensemble is known, we can apply the results of this section in this case, similarly to as we do in Section˜5. The setting that we consider is analogous to the setting that we describe in Section˜1.1.
For a sequence $\{\mu_{N}\}_{N\geq 1}$ of probability measures such that $\mu_{N}\in\mathcal{M}_{N}$ for $N\geq 1$, we let the random variable $p_{k}^{N,\theta_{N}}$ be
$p_{k}^{N,\theta_{N}}\triangleq\frac{1}{N}\sum_{i=1}^{N}\left(\frac{a_{i}}{N\theta_{N}}\right)^{k},$
where $(a_{1},\ldots,a_{N})\sim\mu_{N}$ for $N,k\geq 1$. Similarly, the moments $\{m_{k}\}_{k\geq 1}$ of $\{\mu_{N}\}_{N\geq 1}$ are given by $m_{k}=\lim_{N\rightarrow\infty}\mathbb{E}[p_{k}^{N,\theta_{N}}]$ for $k\geq 1$.
Definition 6.12.
Suppose $\{\theta_{N}\}_{N\geq 1}$ is a sequence of nonzero complex numbers. A sequence $\{\mu_{N}\}_{N\geq 1}$ of probability measures such that $\mu_{N}\in\mathcal{M}_{N}$ for $N\geq 1$ satisfies a Law of Large Numbers with scaling sequence $\{\theta_{N}\}_{N\geq 1}$ and moments $\{m_{k}\}_{k\geq 1}$ if
$\lim_{N\rightarrow\infty}\mathbb{E}_{(a_{1},\ldots,a_{N})\sim\mu_{N}}\left(\prod_{i=1}^{s}p_{k_{i}}^{N,\theta_{N}}\right)=\prod_{i=1}^{s}m_{k_{i}}$
for all positive integers $s$ and $k_{i}$, $1\leq i\leq s$.
The following result is an analogue of Theorem˜1.8, and implies Theorem˜1.8 when $\theta\not=0$.
Theorem 6.13.
Let $\{\theta_{N}\}_{N\geq 1}$ be a sequence of nonzero complex numbers with nonnegative real part such that $\lim_{N\rightarrow\infty}|N\theta_{N}|=\infty$. Let $\{\mu_{N}\}_{N\geq 1}$ be a sequence of probability measures such that for all $N\geq 1$, $\mu_{N}$ is in $\mathcal{M}_{N}$ and is exponentially decaying. For $N\geq 1$ and $\nu\in P^{+}$, define
$c_{\nu}^{N}\triangleq\frac{P(\nu)}{|\nu|!\theta_{N}N}\cdot\frac{\partial}{\partial x_{i_{1}}}\cdots\frac{\partial}{\partial x_{i_{r}}}\ln(G_{\theta_{N}}(x_{1},\ldots,x_{N};\mu_{N}))\bigg{|}_{x_{i}=0,1\leq i\leq N}$
Assume that for all $\nu\in P^{+}$, $|c_{\nu}^{N}|=N^{o_{N}(1)}$ and $|c_{\nu}^{N}|=|\theta_{N}N|^{o_{N}(1)}$. Then, $\{\mu_{N}\}_{N\geq 1}$ satisfies a LLN with scaling sequence $\{\theta_{N}\}_{N\geq 1}$ and free cumulants $\{c_{k}\}_{k\geq 1}$ if and only if
$\lim_{N\rightarrow\infty}\sum_{\nu\in P,|\nu|=k}(-1)^{\ell(\nu)-1}\frac{|\nu|P(\nu)}{\ell(\nu)}c_{\nu}^{N}=c_{k}$
for all $k\geq 1$. Recall that if the free cumulants are $\{c_{k}\}_{k\geq 1}$, then the moments are $m_{k}=\sum_{\pi\in NC(k)}\prod_{B\in\pi}c_{|B|}$ for $k\geq 1$.
Proof.
We follow the proof of Theorem˜1.8 in Section˜6.1; in particular, we use the notation from the proof. Instead of (24), we get that
$=\sum_{p\in\mathfrak{P}_{k}^{N}}s_{p,\lambda}(\theta_{N})N^{k-\ell(p)+\ell(\lambda)}p+O(|\theta_{N}N|^{k}N^{\ell(\lambda)-1+o_{N}(1)})+O(|\theta_{N}N|^{k-1+o_{N}(1)}N^{\ell(\lambda)}).$
Define the complex number $c_{p,\lambda}\triangleq\frac{s_{p,\lambda}(\theta)}{\theta^{k-\ell(p)}}$ for $p\in\mathfrak{P}_{k}^{N}$ or equivalently, $c_{p,\lambda}\triangleq s_{p,\lambda}(1)$; recall that $s_{p,\lambda}(\theta)$ is a complex number multiplied by $\theta^{k-\ell(p)}$, see Lemma˜6.4. Then,
$=\sum_{p\in\mathfrak{P}_{k}^{N}}c_{p,\lambda}(\theta_{N}N)^{k-\ell(p)}N^{\ell(\lambda)}p+O(|\theta_{N}N|^{k}N^{\ell(\lambda)-1+o_{N}(1)})+O(|\theta_{N}N|^{k-1+o_{N}(1)}N^{\ell(\lambda)}).$
Hence,
$\frac{1}{N^{\ell(\lambda)}(\theta_{N}N)^{k}}\sum_{\mathfrak{r}\in I_{N}(\lambda)}[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))$
$=\sum_{p\in\mathfrak{P}_{k}^{N}}c_{p,\lambda}\frac{p}{(\theta_{N}N)^{\ell(p)}}+O(N^{-1+o_{N}(1)})+O(|\theta_{N}N|^{-1+o_{N}(1)}).$
After observing that $c_{\tilde{F}_{N}}^{\nu}=\theta_{N}Nc_{N}^{\nu}$ and using Lemma˜6.3 with $x=1$, we get that
$\frac{1}{N^{\ell(\lambda)}(\theta_{N}N)^{k}}\sum_{\mathfrak{r}\in I_{N}(\lambda)}[1]\mathcal{D}_{\mathfrak{r}}^{\theta_{N}}(\tilde{F}_{N}(x_{1},\ldots,x_{N}))$
$=\prod_{i=1}^{\ell(\lambda)}\left(\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}\left(\sum_{\nu\in P,|\nu|=|B|}(-1)^{\ell(\nu)-1}\frac{|\nu|P(\nu)}{\ell(\nu)}c_{N}^{\nu}\right)\right)+o_{N}(1).$
Afterwards, we can repeat the proof of Theorem˜1.8. ∎